I have 7 CDs I want to select 4 to play in my car CD player. In how many ways can I select 4 of 7 CDs?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

7*6*5*4 = 840
But, selecting A, B, C & D is the same as D,B,A & C
4*3*2*1 = 24
--> 840/24 = 35 ways

Answer by stanbon(75887)

I have 7 CDs I want to select 4 to play in my car CD player. In how many ways can I select 4 of 7 CDs?
----------
You can choose 4 to play or 3 to not play::
Ans: 7C4 = 7C3 = (7*6*5)/(1*2*3) = 35 ways
---------------------
Cheers,
Stan H.
----------------------